OpenStudy (anonymous):

plzzzz helppp: what is a positive value of tan1/2x when sin x=0.8?

OpenStudy (anonymous):

@mossyfish plz helpp

OpenStudy (anonymous):

any1 plz help if u no

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

\[sinx=0.8\] \[cosx=\pm \sqrt{(1-0.8^2)}=\pm0.6\] \[\tan(x/2)=1-cosx)/sinx\] Can you figure this out from here?

OpenStudy (anonymous):

r u using the quadratic formula??

OpenStudy (anonymous):

@mossyfish

OpenStudy (anonymous):

Nope, trigonometric properties. What are your answer choices, btw?

OpenStudy (anonymous):

what is the trigonometric property? canu write it?

OpenStudy (anonymous):

choic are: 0.5 0.4 0.33 0.25

myininaya (myininaya):

if sin(x)=8/10=4/5 then we can find cos(x) by using the right triangle associated with 4/5. Recall \[\cos^2(x/2)=\frac{1}{2}(1+\cos(x)) \text{ and } \sin^2(x/2)=\frac{1}{2}(1-\cos(x))\] so \[\tan(x/2)=\sqrt{\frac{1+\cos(x)}{1-\cos(x)}}\] so basically you are given sin(x)=4/5 just find what cos is off of your right triangle you drew for sin(x)=4/5

OpenStudy (anonymous):

how did u gt sin(x)=8/10=4/5??

myininaya (myininaya):

that is what is given

myininaya (myininaya):

you are given sin(x)=8/10

OpenStudy (anonymous):

no its n the problem is: what is a positive value of tan1/2x when sin x=0.8?

myininaya (myininaya):

right sin(x)=8/10

OpenStudy (anonymous):

0.8 is same at 8/10?

myininaya (myininaya):

of course

OpenStudy (anonymous):

its nt 4/5?

myininaya (myininaya):

8/10=4/5

myininaya (myininaya):

it is called reducing a fraction

myininaya (myininaya):

.8=4/5

OpenStudy (anonymous):

ohh I see

OpenStudy (anonymous):

ok now I don't gt where ur gttin the rest frm?

myininaya (myininaya):

trig identities

myininaya (myininaya):

http://www.purplemath.com/modules/idents.htm

OpenStudy (anonymous):

oh now im gona use these identities to solve the problem?

myininaya (myininaya):

half angle identities

OpenStudy (anonymous):

ok how do I know which identity im gona use for this problem?

myininaya (myininaya):

well the key is noticing it is asking for tan(x/2) that looks like half an angle there because of the over 2 part so we know to use the half angle identity

OpenStudy (anonymous):

so we gona use the tan(x/2)=1+cos x/1-cos x?

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

no you can use \[\tan(x/2)=\sqrt{\frac{1-\cos(x)}{1+\cos(x)}}\]

OpenStudy (anonymous):

so now I need to plug in?

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

after you find the value for cos

myininaya (myininaya):

you know sin(x)=4/5

myininaya (myininaya):

so what is cos(x)=

OpenStudy (anonymous):

ugh can we plz go thro every step? how did we find the value of cos?

myininaya (myininaya):

try to draw a right triangle for sin(x)=4/5 then find the adjacent side brb

OpenStudy (anonymous):

ok waitng

OpenStudy (anonymous):

@mossyfish

OpenStudy (anonymous):

I think @myininaya is doing a fine job helping you :) Tag me in other questions you might need help with.

OpenStudy (anonymous):

she is but she said brb and I thought maybr u could help me for now...

OpenStudy (anonymous):

ik shes helping me good

OpenStudy (anonymous):

Okay well, continuing from where I was from.. When cosx=-0.6 tan(x/2)=(1+06)/0.8-2 When cosx=.6 tan(x/2)=(1-0.6)/0.8 which equals...?

OpenStudy (anonymous):

u both are doin 2 diff things and its confusing me I need u to plz tel step by step if u culd?

OpenStudy (anonymous):

Yeah this is why @myininaya should help you... We both have different methods of solving this, jumping back and forth between us isn't going to help you :( Good luck!

OpenStudy (anonymous):

ok fine thnks

myininaya (myininaya):

Have you drew a right triangle and use the sin(x)=4/5 to label the triangle yet?

OpenStudy (anonymous):

yes

myininaya (myininaya):

then find the missing side

myininaya (myininaya):

which is what?

OpenStudy (anonymous):

do I use the pathagreom theorem?

myininaya (myininaya):

 yep it is a right triangle so use that theorem